OpenStudy (anonymous):

write the 35/6 in improper fraction as a mixed number in simplest form.

OpenStudy (anonymous):

5 5\6

OpenStudy (pixiedust1):

How many times can 6 go into 35?

OpenStudy (unklerhaukus):

35/6 = (5*6+5)/6 = 5+5/6